OpenStudy (anonymous):

A cube is painted red on all six sides, then cut into 125 smaller congruent cubes. A small cube is selected at random and rolled. What is the probability that it has a red side facing up?

OpenStudy (anonymous):

I got 0.176

OpenStudy (anonymous):

I got ------ 8 pieces of cube with 3 sides painted red 36 pieces of cube with two sides painted red and 36 pieces of cube with one side painted red

OpenStudy (dumbcow):

not quite, there are 54 cubes with 1 red side corners have 3 red sides edges have 2 red sides outer faces have 1 red side inner cube have 0 red sides  to get probability, multiply individual probabilities for each case, then take sum probability of getting a corner is 8/125.....chance of landing on red is 3/6 --> (8/125)*(3/6) = 4/125 probability of getting an edge is 36/125.....chance of landing on red is 2/6 --> (36/125)*(2/6) = 12/125 probability of getting outer face is 54/125.....chance of landing on red is 1/6 --> (54/125)*(1/6) = 9/125 of course the probability for inner cubes is zero since they have 0 red sides --> 4/125 + 12/125 + 9/125 = 25/125 = 1/5